OpenStudy (anonymous):

is 2/3 greater equal or less than 5/8???HELP GOLD MEDEL AND FANNNNNNNNNNNN!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (eric_d):

more than

OpenStudy (eric_d):

because

OpenStudy (anonymous):

so 2/3 is greater than 5/8?

OpenStudy (kirbykirby):

You can put everything on the same denominator: \[ \frac{2}{3}=\frac{2\times 8}{3\times 8}=\frac{16}{24}\] \[ \frac{5}{8}=\frac{5\times 3}{8 \times 3}=\frac{15}{24}\]

OpenStudy (eric_d):

Let's wait for @kirbykirby expanation

OpenStudy (anonymous):

can you help me with two more oplease??

OpenStudy (kirbykirby):

now you can easily compare the numerators.. and the largest numerator indicates the overall larger fraction :)

OpenStudy (anonymous):

is 6/3 greater less or equal two 10/15 and write 12/32 in simplset form

OpenStudy (kirbykirby):

Hint: put 6/3 on the same denominator as 10/15 That is... multiply the top and bottom of 6/3 by 5

OpenStudy (texaschic101):

you could also turn the fractions to decimals and compare them that way

OpenStudy (anonymous):

so 6/3 is greater (>)

OpenStudy (kirbykirby):

yep

OpenStudy (anonymous):

nad wite 12/32 in simpleaet form how do you even dot that?

OpenStudy (eric_d):

Divide both numerator and denominator with 4

OpenStudy (texaschic101):

find a highest common number that goes into both numbers....that number would be 4 now divide both numbers by 4 and you get 3/8

OpenStudy (anonymous):

and 2/3>5/8 right

OpenStudy (kirbykirby):

yes..

OpenStudy (texaschic101):

first find common denominator for both numbers...that would be 24. now convert the fractions where they have the denominator of 24 2/3 = 16/24 5/8 = 15/24 so .... 2/3 > 5/8 or 5/8 < 2/3

OpenStudy (anonymous):

how do you wite 3 and 2/3 as a improper fraction????

OpenStudy (texaschic101):

take the whole number...multiply it by the denominator, then add that to the numerator...and put that over the original denominator

OpenStudy (anonymous):

how?/ can youm draw it

OpenStudy (texaschic101):

3 2/3.....take the whole number (3)...multiply it by the denominator (3) and you get 9...now add it to the numerator(2) and you get 11...now put that over the original denominator (3) and you end up with 11/3

OpenStudy (anonymous):

last on and i will do the rest by my self

OpenStudy (texaschic101):

ok

OpenStudy (anonymous):

so wite 17/4 as amixed number

OpenStudy (anonymous):

ohh and I just gave you medel and I will fan you it just a bit

OpenStudy (anonymous):

ok just fanned u

OpenStudy (texaschic101):

how many times does 4 go into 17 without going over ?.It goes in there 4 times. So 4 is your whole number. and since 4 x 4 = 16, that leaves you with a remainder of 1, this is your numerator....and you use the same denominator, which is 4. So 17/4 in a mixed number = 4 1/4

OpenStudy (anonymous):

thank yuo sooooooooooooo much i just fanned you and gave you a medel!!! ;)

OpenStudy (texaschic101):

sweet...glad to help :)

OpenStudy (anonymous):

yup

OpenStudy (texaschic101):

if you have any more problems, just tag me and I will be happy to help

OpenStudy (anonymous):

ohh and just to be sure 6/3>10/15 rigtht

OpenStudy (texaschic101):

6/3 = 30/15 10/15 6/3 > 10/15....yep, your right

OpenStudy (anonymous):

ok and 2/3 > 5/8

OpenStudy (anonymous):

phew I jst want to say A BIGGGGGG thank you for all you time and you help

OpenStudy (texaschic101):

2/3 = 16/24 5/8 = 15/24 2/3 > 5/8...yep I am glad to help...thats what I am here for :)

OpenStudy (anonymous):

:D

OpenStudy (texaschic101):

good luck with your math :) any problems, just tag me